# Retrieve, Summarize, Plan: Advancing Multi-hop Question Answering with an Iterative Approach

### v1

@@ 5.3 Case Study
In the first case, the retrieved documents are consistent since the initial retrieval question is the same. However, the two models make different decisions about what to do next. IRCoT, which integrates information processing and planning in one step, has a higher level of task complexity. It mistakenly misses information related to "Twisted Fortune", which leads the model to decide that further retrieval is needed, resulting in over-planning. On the other hand, ReSP accurately and comprehensively acquires the supporting facts related to the overarching question through the summarizer. Consequently, the reasoner determines that the question can be answered, and the generator produces the correct response, thereby halting the iteration after the first round of retrieval.
In the second case, at the end of the first round of retrieval, both models make similar decisions due to the absence of information related to the main subject, "Rachelle Amy Beinart" in the documents: they both query for documents related to "Rachelle Amy Beinart". However, due to limitations in document coverage or retriever capability, no relevant documents on "Rachelle Amy Beinart" are found. At this point, the two models propose different sub-questions. Lacking a recorded retrieval trajectory, IRCoT can only make judgments based on the current information, thus continuing to query "Rachelle Amy Beinart", leading to repetitive planning. In contrast, ReSP, which avoids outputting previously retrieved sub-questions, adjusts the retrieval subject and produces a sub-question about "film that follows a group of rebels on a mission", thereby avoiding repetitive planning.
6 Conclusion
@@
Appendix A Prompt Templates of Modules
The prompt templates of different modules in ReSP are shown in Table A.
Module Function Prompt
[…]
Table 4: The prompt templates of different modules in ReSP. The input parameters include Overarching question: the input question; Sub-question: the current sub-question generated by the reasoner; Combined memory queues: the concatenated content of the global evidence memory and local pathway memory queues for each iteration; docs: the documents retrieved from a single round of retrieval.

### v2

@@ 6.5. Case Study
In the first case, the retrieved documents remain consistent because the initial retrieval question is the same. Despite this consistency, the two models take different approaches to their subsequent actions. IRCoT, which combines information processing and planning in a single step, tackles a higher level of task complexity. Unfortunately, it fails to capture information related to "Twisted Fortune", which leads the model to conclude that additional retrieval is necessary, thereby leading to over-planning and unnecessary steps. In contrast, ReSP effectively and comprehensively gathers the necessary supporting facts related to the overarching question by utilizing the summarizer. As a result, the reasoner accurately assesses that the question can be answered with the information at hand. Consequently, the generator can produce the correct response, allowing the process to conclude successfully after just the first round of retrieval, without the need for further iterations.
In the second case, after the first round of retrieval, both models arrive at similar decisions due to the absence of information concerning the main subject, "Rachelle Amy Beinart" in the retrieved documents. As a result, both models initiate queries for documents specifically related to "Rachelle Amy Beinart". However, limitations in document coverage or the capabilities of the retriever result in no relevant documents being found on "Rachelle Amy Beinart". At this juncture, the models adopt different approaches in formulating sub-questions. Without a recorded retrieval trajectory, IRCoT is constrained to make decisions solely based on the current set of information, which leads it to continue querying "Rachelle Amy Beinart", resulting in repetitive and unproductive planning. In contrast, ReSP employs a strategy that avoids generating previously retrieved sub-questions. This allows it to adapt by adjusting the retrieval subject and crafting a new sub-question focused on "film that follows a group of rebels on a mission". This shift in focus helps ReSP circumvent repetitive planning, enhancing the efficiency and effectiveness of the retrieval process.
7. Conclusion